OpenStudy (megans20):

I need help studying for the Algebra 1 EOC please.

OpenStudy (andrejrusa12):

@hartnn @TheSmartOne @Jadeishere @welshfella @Nnesha @JoannaBlackwelder @ganeshie8 @JFraser @Loser66 @Vuriffy @skullpatrol @Apple @ILovePuppiesLol @imqwerty @AloneS @Ambuj @agent0smith @jhonyy9 @mathstudent55 @whpalmer4 @Wolf @Ellie201 @rayep12 @rosielopez_17 @Revircs @rbrasch @Tifferslovesyaoi @TinaSh @Ttrichards @todjunior @ttttttt @YanaSidlinskiy @yasatg @UnkleRhaukus @Unavailabilityy @userk2 @oliviad @oopsydaisy35 @poopsiedoodle @PandaNinja101 @procrastilate @pooja195 @Agl202 @Aleah54 @AAbomosalam1998 @Autumn_12001 @Ashton_Bishop @anastasiai @sweetburger @samanthagreer @something @sam_OVO @smollberru @sophiesky @Daniellelovee @Dyann.Cardona @ddoll01 @Deadpool @Finette @gudetamas @Goldendragon01 @Hayhayz @HocusPucus292333393 @hadeel @htxatlarge @jtug6 @johnsmith31613 @jjjjjjjjjjjjhujljnjb @Jaynator495 @jim_thompson5910 @jabez177 @kewlgeek555 @kidrah69 @kellyspeakslouder @k_d_1 @kitten_dude @leenathan @Lacris @lyss @ljetibo @Lyssburleson @Loli @zzr0ck3r @zepdrix @basically @breanna.rae @Batman4ever @nerdkid @noma @MarcoReus11 @mantar0078 @Makayla_Caldwell @kevin @will.h

Arcadiouse (arcadiouse):

Things I would go over is, inequalities, equations

OpenStudy (megans20):

I'm taking the practice test now, but I took Algebra 1 three years ago and I don't remember any of this stuff

Arcadiouse (arcadiouse):

Well, whatever you fail on the practice test study. That is what its for.

OpenStudy (megans20):

I'm not sure how to study for it, I tried to look up one problem but I can't find anything close to it. Could you help me with the problem I'm stuck on?

Arcadiouse (arcadiouse):

Maybe.

OpenStudy (mathgeek289):

Yup

OpenStudy (megans20):

Thanks. I'll post the question. I just need guided through it, or instructions on what to do for it.

OpenStudy (megans20):

The table below shows the values for the function y=f(x). x: -4, -1, 0, 3, 6 y: 7, -2, 3, -4, 5 Complete the table for the function y=f(1/5x). Then it gives me a blank table.

OpenStudy (megans20):

Sorry it it's messy, I wasn't able to copy it from the practice test so I just had to type everything

OpenStudy (mathgeek289):

@Arcadiouse

zepdrix (zepdrix):

Are you ok with function notation? Like if I write this,\[\large\rm f\left(\color{orangered}{-4}\right)=7\]does it make sense to you? I got that information from the first chart. `-4 in the input` and `7 is the output`.

OpenStudy (megans20):

I wouldn't say I'm an expert on function notation, but that doesn't look too unfamiliar to me

zepdrix (zepdrix):

Notice with this new input,\[\large\rm f\left(\color{orangered}{\frac15x}\right)=?\]If we chose to put x=-4 like before,\[\large\rm f\left(\color{orangered}{-\frac{4}{5}}\right)=?\]It gives us a problem. We don't know enough about the function to determine what the output would be when the input is -4/5.

zepdrix (zepdrix):

But what we do know... is that the output is 7 when the input is -4. So what if we try something sneaky. What if we try to change this entire input \(\large\rm \color{orangered}{\frac{1}{5}x}\) into \(\large\rm \color{orangered}{-4}\)? Then we would be able to determine that we have 7 again, right?\[\large\rm f\left(\color{orangered}{\frac15x}\right)=f(\color{orangered}{-4})=7\]

zepdrix (zepdrix):

Maybe I can put this in simpler words.. umm We're going to take the old x values, set them equal to this new input, and solve for the new x values.\[\large\rm \frac15x=-4\]Multiplying by 5 gives us our first `new x value`, \(\large\rm x=-20\) But our y values will all be the same.

zepdrix (zepdrix):